OpenStudy (nerdrage29):

HELP ME SOLVE THIS. sin(2 theta)-cos(theta)-2 sin(theta)+1=0

OpenStudy (nerdrage29):

I know you can use the double angle formula and whatnot, but im still baffled. Also, it asks to find the values only in 0 to 2pi.

OpenStudy (nerdrage29):

please don't give the answer. Just show me clues to solving this.

OpenStudy (nerdrage29):

What would you recommend?

OpenStudy (nerdrage29):

This is precalculus.

OpenStudy (zpupster):

use the trig identity sin(2 theta)=2sin(theta)cos(theta)

OpenStudy (nerdrage29):

i did

OpenStudy (nerdrage29):

im still stumped. mothing can be factored after

OpenStudy (agent0smith):

\[\large 2\sin \theta \cos \theta -\cos \theta-2 \sin \theta +1=0\] from wolfram looks like you have to use the half angle formulas to solve this... see "alternate forms" http://www.wolframalpha.com/input/?i=sin(2+theta)-cos(theta)-2+sin(theta)%2B1%3D0

OpenStudy (nerdrage29):

You used wolfram to solve it?

OpenStudy (nerdrage29):

How may I use wolfram to study better?

OpenStudy (nerdrage29):

I tried to use it but there were so many different calcluators.

OpenStudy (agent0smith):

No i'm saying look at the alternate forms, as a way to help solve it. It is not at all obvious, though.

OpenStudy (nerdrage29):

I also want to see the steps wolfram used to solve this. All I get is the final outcome. I want to learn how to derive it.

OpenStudy (agent0smith):

Hey look...\[\large 2\sin \theta \cos \theta -\cos \theta-2 \sin \theta +1=0\] \[\large \cos \theta(2\sin \theta -1)-1(2 \sin \theta -1)=0\]

OpenStudy (agent0smith):

\[\large (\cos \theta-1)(2\sin \theta -1)=0 \]

OpenStudy (nerdrage29):

My goodness. You are brilliant.

OpenStudy (nerdrage29):

I got to your first step, but I didn't picture the negative 1 as being a factor..

OpenStudy (agent0smith):

Haha thank you

OpenStudy (nerdrage29):

How did you think of that?

OpenStudy (agent0smith):

Tried factoring cosine out of the first two terms, noticed the similarity with the remaining terms

OpenStudy (agent0smith):

Sometimes trial and error pays off.

OpenStudy (nerdrage29):

2sin(theta)cos(theta)-cos(theta)-2sin(theta)+1 Then, Cos(theta)*(2sin(theta)-1)-2sin(theta)+1 Then, (cos(theta) +1) (2sin(theta)-1) = 0 For my solution, I got pi, pi/6/ and 5pi/6. However my friend said it was 0 instead of pi. Can someone show my error?

OpenStudy (nerdrage29):

Ahh, i distributed it backwards, and found my error. Thank you.